3. Consider the problem regarding the information forms about survival rates of surgery and radiation treatments from class. Suppose the number of surgeons who receive the forms in the two groups are RA 100 Let the population proportions of doctors receiving the forms whod recommend surgery be denoted PA and p (a) Formulate a null and alternative hypothesis to investigate the possibility that surgeons are affected by the language used to convey survival rates (b) Using type I error rate a.05, define a (large-sample) rejection region for the test statistic given below: (c) Suppose the observed data ar A0.5 and B0.63. What is your decision regarding Ho? (d) Repeat part (b) using a10 (e) Repeat part (c) using a 10. (f) Using large-sample theory, what is the smallest level a you could have used bs sill nijected Ho with the data froams part (
Answer #1

bu the lanuage 4: Ph-k ο ( Surgeons are not affected by the lan nage wed) イ o.os一 +もん: t 1.9 사 A冫0.63 几 nA

f) The smallest level of alpha for which we could reject H0 is a p-value.

p-value = 2(1- \phi (|Z0|) = 2(1-\phi(|1.87|) = 2(1-0.9696) = 0.06148.

the minimum level of aplha = 0.06148.